OpenStudy (annie96):

1.) Write 6x-3y=3 in slope form. Show all work! 2.) Based on your answer to the question above, state the slope and the y-intercept of the line. 3.) Based on the two previous questions. Find the line that is perpendicular to this line and passes through (4,-2).

OpenStudy (texaschic101):

1. 6x - 3y = 3 (subtract 6x from both sides) -3y = -6x + 3 (divide by -3) y = -6/-3x + 3/-3 y = 2x - 1 <== slope intercept form 2. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. Therefore, the slope of this line is 2 and the y intercept is -1. 3. If you are looking for a perpendicular line, you need the negative reciprocal of the slope given. All that means is " flip " the slope and change the sign. The slope given is 2 or 2/1, the negative reciprocal slope you need is -1/2 (see how I flipped the slope and changed the sign... Now we will use y = mx + b and solve for b because we already know that the slope is -1/2. y = mx + b slope(m) = -1/2 (4,-2) x = 4 and y = -2 now we sub y = mx + b -2 = -1/2(4) + b -2 = - 2 + b -2 + 2 = b 0 = b your perpendicular line is : y = -1/2x + 0

OpenStudy (annie96):

Thank you.

OpenStudy (texaschic101):

no problem :)